Find the marginal cost functions for each of the following average cost functions. a) AC = 1.5Q + 4 + 46 𝑄 b) AC = 160 𝑄 + 5 - 3Q + 2Q2


1
Expert's answer
2022-03-16T11:27:48-0400

Solutions

 "MC=\frac{dTC}{dQ}"


"TC = AC \times Q"


"a)AC = 1.5Q + 4 + 46Q"


Differentiate the AC function


"TC=1.5Q^{2}+4Q+46Q^{2}"


"MC=3Q+4+98Q"


"MC=101Q+4"



"b)AC=160Q+5-3Q+2Q^{2}"


"MC=\frac{dTC}{dQ}"


"TC=AC\times Q"


"TC=160Q^{2}+5Q-3Q^{2}+2Q^{3}"


"MC=320Q+5-6Q+6Q^{2}"


"MC=314Q+5+6Q^{2}"